Marks of six students are tabulated below:

Name : Raj Anil Amit Om Rita Renu

Marks : 54 50 52 48 50 52

From the population, tabulated above, you are supposed to choose a sample of size two.

(a) Determine, how many samples of size two are possible

(b) Construct sampling distribution of means by taking samples of size 2 and organize the data.




1
Expert's answer
2021-05-13T07:09:04-0400

(a)

"Number = \frac{N!}{n!(N-n)!} \\

n = 2 \\

Number = \frac{6!}{2!(6-2)!} = 15"

(b)